College softball distance from the pitcher's mound to the batter is 43 feet. If the ball leaves the bat at 85 mph, how much time elapses between the hit and the ball reaching the pitcher?
1
Expert's answer
2012-05-25T10:15:36-0400
1 Mile = 5280 Feet
So, the time which elapses between the hit and the ball
reaching the pitcher:
t = d / v = 43 / (85 * 5280) = 43 / 448800 hrs
To
get the answer in seconds, we should multiply it by 60 * 60 = 3600.
t = 3600
* 43 / 448800 = 0.344919786 = approx. 0.345 seconds.